OpenStudy (anonymous):

help pls.... using leibnitz's theorm find nth differential coefficient of cos2x.x^2

OpenStudy (anonymous):

obtained this step

OpenStudy (anonymous):

but the final answer is :

OpenStudy (anonymous):

well lets start by breaking it down

OpenStudy (anonymous):

Let's see if that second formula is correct. I would say it's not because it doesn't account for the \(2x\) within the cosine. It's more likely the formula for the \(n\)th derivative of \(x^2\cos x\). \[\begin{align*} f(x)&=x^2\cos2x\\\\ f'(x)&=2x\cos2x-2x^2\sin2x \end{align*}\] The given answer says \[f^{(n)}(x)=x^2\cos\left(x+\frac{n\pi}{2}\right)+2nx\cos\left(x+\frac{(n-1)\pi}{2}\right)\\ \quad\quad\quad\quad+n(n-1)\cos\left(x+\frac{(n-2)\pi}{2}\right)\] which would imply \[f'(x)=x^2\cos\left(x+\frac{\pi}{2}\right)+2x\cos(x-\pi)=x^2\sin x-2x\cos x\] which is not the same as the actual computed derivative.

OpenStudy (anonymous):

thank you!!.. i confused with the question. As you said, the second one is the final answer for x^2 cosx and the first one is for x^2cos2x